The two PUPPIs that were deployed at Dead Dog mound on 29 August were recalled and recovered. Initial analysis of the pressures recorded by the PUPPIs indicates that the shallow sediment is underpressured relative to the seafloor (i.e., the area is an inflow zone).

Hole 1035F (Prospectus Site BH-7)

The ship was moved in DP mode back to Hole 1035B and offset 10 m south to improve the 3-dimensional image of the Bent Hill massive sulfide deposit by using rotary coring to recover the deeper sections. The hole was at the base of the Bent Hill massive sulfide mound and 40 m south of Hole 856H. The same RCB BHA was run with an MBR to release the bit in case the pipe became stuck at the bit. The seafloor was assumed to be at 2459.0 mbrf. Hole 1035F was spudded at 1500 hr on 17 September. RCB Cores 169-1035F-1R to 23R were taken from 0 to 224.8 m (2683.8 mbrf) with 10.0% recovery. The upper hole was unstable, and two sepiolite mud sweeps were required to stabilize the hole. Torque and hole fill were controlled with sepiolite mud sweeps every other core. Five of 23 cores were jammed in the bit throat or core catcher. The plan was to return to Hole 1035F for logs after completing other work; therefore, a Free Fall Funnel (FFF) was deployed. The FFF position was concealed by heavy flow coming out of the hole; however, the FFF appeared to be in position based on sonar returns. The bit cleared the seafloor at 1000 hr.

Hole 1035G (Prospectus Site BH-3)

The ship was moved in DP mode back to Hole 1035A and offset 10 m east (near the base of the mound) to recover the deeper sections. The hole was 65 m west of Hole 856H. The seafloor was tagged at 2456.0 mbrf while the TV was down. Hole 1035G was spudded at 1245 hr on 19 September. The hole was washed from 0 to 44.4 m with a center bit through the silty clay interval, and wash barrel Core 169-1035G-1W had no recovery. RCB Cores 169-1035G-2R to 5R were taken from 44.4 to 83.3 m in the massive sulfide and stringer zone. The hole was washed from 83.3-140.9 m in silty claystone, and wash barrel Core 169-1035G-6W had 4.5 m recovery. RCB Cores 169-1035G-7R to 13R were taken from 140.9 to 208.5 m (2664.5 mbrf). Coring was terminated at the target depth. The VIT-TV was run to the seafloor, and a 15 in. hole was observed with a small cuttings mound, no crater, and no visible flow in or out.

Return to Hole 1035F

The ship was moved in DP mode back to Hole 1035F with the VIT-TV down. A strong flow was observed coming out of the FFF throat and billowing upward to more than 40 m above seafloor with abundant suspended debris.

Hole 1035H (Proposed Site BH-8)

The ship was moved in DP mode to the mound 200 m south of the BHMS deposit (now called ODPMS mound). The mound was surveyed for a good core site on the upper plateau. Hole 1035H was spudded at 2030 hr on 20 September on a reasonably flat bench, adjacent to an inactive toppled chimney and 8 m from the top of the mound. The seafloor was tagged at 2455 mbrf. RCB Cores 169-1035H-1R to 16R were taken from 0 to 142.3 m (2597.3 mbrf) with 12.5% recovery. The standard butyrate (tenite) core liners are rated to 150°C service. On Core 169-1035H-11R, the butyrate liner melted completely (into a white lump around the core) during a 20-minute shutdown when the overshot sheared. The butyrate core liners on Cores 169-1035H-12R to 14R were partially melted.The quick-release polypak seal at the top of the RCB inner barrel (OD3203) was melted on some runs.

After Core 169-1035H-16R at 142.3 m, the pipe was pulled up to 74.3 m, and the Los Alamos water sampler was run. The sampler timer slid back in the case and failed to go off. A temperature tab on the sampler turned black out to 399°C. The VIT-TV was run to the seafloor, and a strong flow was observed coming from the hole. RCB Cores 169-1035H-17R to 27R were taken from 142.3 to 247.9 m with 11.1% recovery. Coring was terminated due to time constraints. The Ultem core liners were used well above the design temperature of 200°C on Cores 169-1035H-15R to 27R by maintaining circulation at 175 gpm. The liners are extrusion fabricated at 330-357°C. On Core 169-1035H-26R, an Ultem liner melted (like shrink wrap) around the core and sucked into a V shape above the core. The core was cold when recovered.

On Core 169-1035H-26R, the torque increased from 150 to 200 amps and pressure increased from 400 to 550 psi. The pipe was pulled to 74.5 m, and a perforated and blanked off core catcher sub was run with various temperature sensitive tabs and materials (lead, tin, and engineering materials such as O-rings and seals). The temperature tabs indicated a 316°C flowing temperature; however, water may have invaded the tab. The butyrate core liner samples were melted and black, the 200°C Ultem liner sample was melted, the APC seal ring was dissolved, but the viton and 70 and 90 durometer O-rings were not affected. The Los Alamos water sampler was run to 74.5 m but failed to get a sample when the O-rings and valve seat failed. The VIT-TV was run to check the wellbore, and sample nets were attached to the VIT frame to trap material being expelled from the hole. There was a heavy and relatively-clear flow out of the hole with small flakes being blown up to 22 m above the hole. The VIT sample net was run to 6 m above the seafloor and recovered anhydrite flakes, sediment pebbles, and massive sulfide flakes to 2-cm diameter. Some sea surface shrimp and crustaceans were also recovered. The bit cleared the rotary at 0530 hr on 23 September.

The drill pipe that had been below the seafloor and the BHA were coated with 1/16 in. of anhydrite and clay. All the connections were broken, checked, and redoped as a precaution. The dope had been dried out, and the threads were dry. The bottom two drill collars had to be broken out with rig tongs; however, there was no obvious thread damage. The MBR was cemented up with anhydrite and had to be junked. The 9-7/8" RBI CC-7 bit was graded T7B7, SF, BT, CT, LT, I, TD, NR, and had extensive tooth loss in the heel rows, some broken and chipped teeth in the intermediate rows, four loose cones with failed seals, and minor stabilizer gage wear. The bit cored 681.2 m in massive sulfides in 38.4 rotating hours, and logged at least 116 hr in hot holes. The beacon was recovered.


Vancouver Island Helicopters chopper arrived at 1130 hr on 22 September to medivac an ODL employee suffering from internal bleeding. The chopper took 75 gal of helifuel and departed at 1200 hr for Victoria General Hospital.

To 169 Site 1037

To 169 Table of Contents

Publications Homepage

ODP Homepage